OpenStudy (anonymous):

Please help I really don't know how to find decomposition. write the partial fraction decomposition of the given rational expression. x^2/ (x-2)^2(x+1)

OpenStudy (anonymous):

is it yours? \[\frac{ x^2 }{ (x-2)^2(x+1) }\]

OpenStudy (anonymous):

take a look at http://www.purplemath.com/modules/partfrac2.htm everything there, just imitate

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

look at page 2 at the link I give, if you still not understand, let me know then

OpenStudy (anonymous):

alright thanks for the link!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I just want to double check if this is the right answer \[\frac{ \frac{ 1 }{ 9 } }{ (x+1) }+\frac{ \frac{ 13 }{ 18 } }{ (x-2) }+\frac{ 1 }{ (x-2)^{2} }\]

OpenStudy (anonymous):

@Hoa

OpenStudy (anonymous):

what'sup?

OpenStudy (anonymous):

I just wanted to double check the answer above @Hoa if its right?

OpenStudy (anonymous):

so the answer is actually what you have? and not what I got? @phi

OpenStudy (phi):

Yes. That is from wolfram. I did not go through the details, but I assume it is correct.

OpenStudy (anonymous):

thank you i will try to figure out how it got to the final answer!